.Solve𝑥 ≡ 7(𝑚𝑜𝑑 11) 𝑥 ≡ 6(𝑚𝑜𝑑 8) 𝑥 ≡ 10(𝑚𝑜𝑑 15) Also find the


smallest non-negative solutions

1
Expert's answer
2021-12-15T06:20:22-0500

Let us solve the following system of congruences:


"\begin{cases}
x≡7 \mod 11\\
x≡6\mod 8\\
x≡10\mod 15
\end{cases}."


The first congruence is equivalent to the equality "x=7+11t,\ t\in\Z." Let us put this in the second conqruence. Then we have "7+11t\equiv 6 \mod 8." The last conqruence is equivalent to "11t\equiv -1 \mod 8," and hence to "(11-8)t\equiv (16-1) \mod 8," that is "3t\equiv 15 \mod 8." Since 3 and 8 are relatively prime, we conclude that "t\equiv 5\mod 8." It follows that "t=5+8s," and hence "x=7+11(5+8s)=62+88s." Let us put this in the last conqruence. Then we get "62+88s≡ 10 \mod 15," which is equivalent to "88s≡-52\mod 15," and hence to "13s≡-52\mod 15." It follows that "s\equiv-4\mod 15," and hence "s=-4+15k." We conclude that "x=62+88s=62+88(-4+15k)=-290+1320k."


Consequently, the solution of the system is "x\equiv -290\mod 1320" or "[-290]_{1320}."


The smallest non-negative solution is "-290+1320=1030."